I really need help:
In 5 years Chim will be three times as old as Jared. The sum of their ages is 42. How old are Chim and Jared.

Answer by awang1996(50)

In 5 years Chim will be three times as old as Jared. The sum of their ages is 42. How old are Chim and Jared.


5(c+5)=j+5
c+j=42
5c+25=j+5
c+j=42
5c+20=j
c+j=42
c+5c+20=42
6c+20=42
6c=22
c=3.666666666666
j=38.3333333333333333333